OpenStudy (anonymous):

√x-2=x-8 Solve: {11} {6, 11} {}

OpenStudy (solomonzelman):

\[\sqrt{x}-2=x-8\] first say let sqrt x=a, so you will now have, a-2=a^2-8 can you solve this?